A lake manager wants to stock a small lake with 11,550 fish, some bass and some walleyes. He wants to use 1830 more bass. How many of each kind of fish does he need?
Found 3 solutions by Alan3354, addingup, MathTherapy:
Answer by Alan3354(69443)

A lake manager wants to stock a small lake with 11,550 fish, some bass and some walleyes. He wants to use 1830 more bass. How many of each kind of fish does he need?
=================
B + W = 11550
B - W = 1830
------------------------------- Add

Answer by addingup(3677)

b + w = 11,500 (1)
b = w + 1,830
~~~~~~~~~~~~~~~~~~~~~~~~
In (1), substitute for b:
w + (w + 1830) = 11,500
2w = 9,670
w = 4,835
He needs 4,835 walleyes and 4,835+1,830 = 6,670 bass

Answer by MathTherapy(10552)


A lake manager wants to stock a small lake with 11,550 fish, some bass and some walleyes. He wants to use 1830 more bass. How many of each kind of fish does he need?
Correct answers: